OpenStudy (anonymous):

1=2-4e^-(x-1)^2 determine values of x I have gotten to: -(x-1)^2=ln(1/4) and now not sure what to do

OpenStudy (anonymous):

\[1=2-4e ^{-(x-1)^2}\]

OpenStudy (mathstudent55):

\(\large 1=2-4e ^{-(x-1)^2}\) Subtract 2 from both sides. Then divide both sides by -4.

OpenStudy (anonymous):

I am at: -(x-1)^2=ln(1/4)

OpenStudy (mathstudent55):

\(\large \dfrac{1}{4}= e ^{-(x-1)^2}\) \(\large \ln\dfrac{1}{4}= \ln e ^{-(x-1)^2}\) \(\large \ln\dfrac{1}{4}= -(x-1)^2\) \(\large -(x-1)^2 = \ln\dfrac{1}{4} \) Good. I got there too. Now multiply both sides by -1. Then take the square root of both sides being careful to set the right side as +/- the square root. Then add 1 to both sides of both equations. as in: \(x^2 = k\), then \(x = \pm \sqrt{k} \)

OpenStudy (mathstudent55):

Also, \(ln \dfrac{1}{4} = \ln 2^{-2} = -2 \ln 2\). Which may be a more elegant way of writing the log part.

OpenStudy (anonymous):

when I multiply be negative 1 does that give me (x-1)^2=??

OpenStudy (mathstudent55):

Yes. \(\large -(x-1)^2 = -2\ln 2\) \(\large (x-1)^2 = 2\ln 2\)

OpenStudy (anonymous):

ahhhhh ok so it could also be (x-1)^2=ln4

OpenStudy (mathstudent55):

Yes.

OpenStudy (mathstudent55):

Now take the square root of both sides, and use the +/- on the right side before the root sign. Then add 1 to both sides.

OpenStudy (anonymous):

\[x=\pm \sqrt{\ln4}+1\]

OpenStudy (mathstudent55):

Correct. \(\large x-1 = \pm \sqrt{2\ln 2}\) \(\large x =1 \pm \sqrt{2\ln 2}\)

OpenStudy (mathstudent55):

Your answer is fine.

OpenStudy (anonymous):

Now I need to find the derivative of \[y=2-4e ^{-(x-1)^{2}}\]

OpenStudy (anonymous):

I have a mental blank

OpenStudy (anonymous):

is it \[-4e ^{-(x-1)^{2}}something?\]

OpenStudy (mathstudent55):

u is a function of x Then: \(\large \dfrac{d}{dx} e^u = e^u u'\)

OpenStudy (mathstudent55):

\(\dfrac{d}{dx} e^x = e^x \cdot \dfrac{d}{dx} x\), but \(\dfrac{d}{dx} x = 1\), so \(\dfrac{d}{dx} e^x = e^x \)

OpenStudy (mathstudent55):

In your case, the exponent is not simply x, so you need to take its derivative too.

OpenStudy (mathstudent55):

\(\large y=\color{red}{2}-4e ^{\color{green}{-(x-1)^{2}}}\) \(\large y'=\color{red}{0}-4e ^{-(x-1)^{2}}\color{green}{(-1)(2)(x - 1)}\) \(\large y'=8(x - 1)e ^{-(x-1)^{2}}\)

OpenStudy (mathstudent55):

I hope the colors make it clearer.

OpenStudy (anonymous):

Thank you so very very much for all of your help, I really appreciate it :)

OpenStudy (mathstudent55):

You're very welcome.